# Conceptwm: A Diffusion Model Watermark for Concept Protection

### v1

@@ Abstract
The personalization techniques of diffusion models succeed in generating specific concepts but also pose threats to copyright protection and illegal use. Model Watermarking is an effective method to prevent the unauthorized use of subject-driven or style-driven image generation, safeguarding concept copyrights. However, under the goal of concept-oriented protection, current watermarking schemes typically add watermarks to all images rather than applying them in a refined manner targeted at specific concepts. Additionally, the personalization techniques of diffusion models can easily remove watermarks. Existing watermarking methods struggle to achieve fine-grained watermark embedding with a few images of specific concept and prevent removal of watermarks through personalized fine-tuning. Therefore, we introduce a novel concept-oriented watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. We conduct extensive experiments and ablation studies to verify our framework. Our code is available at https://anonymous.4open.science/r/Conceptwm-4EB3/.
1 Introduction
@@
Recent advancements in diffusion models have revolutionized the field of image generation, making it possible to create realistic images from simple text description inputs. With the growing demand for generating highly personalized and customized content, the personalization techniques [29, 11, 10, 18] of diffusion models have achieved significant progress in generating novel concept descriptions (e.g., unique visual themes, motifs, stylistic preferences, or characters) with a few images of specific themes or styles. Through powerful personalization tools like DreamBooth [27] and Lora [14], anyone, even without design skills, can create images of specific concepts using just a few personal images and simple text inputs.
@@
Figure 1: (a) Post-processing Watermark. (b) U-net Finetuning Watermark. (c) VAE-Finetuning Watermark. (d) Latent-based Watermark. (e) Ours.
However, the use of personalization techniques introduces significant risks, particularly the potential for concept misuse or leakage. Untrustworthy users can exploit these tools to create unauthorized concepts related to celebrities or artworks, leading to deepfake incidents [15], the spread of fake news and online spam [4], and serious copyright violations, as the original creators’ intellectual property and creativity are undermined.
@@
To address the ownership and accountability issues of specific concepts in generative models, recent research has introduced the task of concept watermarking [7], which aims to embed watermark information to the concept and extract the watermark from the subsequent generated images based on the watermarked concept. In practice, the model owner typically possesses a pre-trained diffusion model and seeks to enable the generation of images for a new concept. The concept owner provides images with a specific concept to fine-tune the existing diffusion model while embedding watermark information related to the concept, enabling the model to generate images containing the new concept and allowing the watermark decoder to extract watermark information from the generated images.
@@
Even though adding watermarks to generated images has brought increasing attention [17, 36, 41, 37, 9, 43, 34, 8, 6], it is non-trivial for existing watermarking solutions to achieving concept watermarking. On one hand, most existing methods [34, 9, 43] typically watermark the whole images rather than images of specific concepts, so that they are inapplicable in the concept watermarking scenario. Fine-grained watermarking mechanisms for specific concepts has rarely addressed by prior work. The major challenge is that imgae watermarking concept watermarks must not only exist within a specific image, but also remain present across various image variations generated by diffusion model. Dissimilar to traditional image watermarks embedded at the pixel level, the concept watermarks appears a higher-level technical barrier as it associates the watermarks with high-levels of sematic features related to specific concepts. Since the amount of images with a specific concept generally is limited for personalization model fine-tuning by injecting a new concept, it is a great challenge for existing methods which embed watermarks with a large volume of training samples.
@@
On the other hand, robustness is another challenge for existing watermarking methods, when considering diverse adversaries in the concept protection scenarios. Recent work also showed that existing methods failed in defending model purification attacks [45, 21], as adversaries can bypass watermarking schemes by generating various watermark-free images through personalized fine-tuning. Balancing adversariality, image quality, and watermark fidelity becomes a technical obstacle. One of alternatives for increasing watermark adversariality is to inject imperceptible protective adversarial perturbations into the protected images, but this idea generally has a negative impact on both image quality and watermark fidelity [19, 28, 31]. Therefore, finding out a watermarking solution to ensuring concept tracing while addressing threats from adversaries has an urgent demand.
@@
There are three core properties needed for satisfying the requirement of concept watermarking, namely, refined nature, robustness, and adversarial properties. Refined nature: The knowledge of a specific concept is learned by the denoising network (U-net), which is the core component of the diffusion model. Adding a watermark to a specific concept should be done with the least amount of training data and training steps possible, while ensuring that the fidelity before and after adding the watermark remains unaffected. Robustness: The watermark should remain stable across various image processing operations and under different generation parameters. Adversarial properties: The generated concept-specific images should not be subject to “jailbreaking” via personalized fine-tuning.
@@
To address the challenges outlined above, we propose a novel concept watermarking framework ConceptWm. We consider that a refined watermark training and Fidelity Preserving Perturbation are two critical factors in the context of concept watermarking. To achieve the concept watermark embedding, we propose Concept-oriented Watermarking Finetuning, which ensures the rapid learning of watermark embedding patterns in diffusion models with limited data and training budget. To ensure the robustness of the watermark against various image processing attacks and facilitate the learning of the watermark diffusion model, our Watermark Components Pretraining Module incorporates the watermark into the latent space and employs a distortion layer to ensure the watermark can withstand various image processing attacks. To effectively balance adversarial perturbation, watermark fidelity, and image quality, while mitigating the risk of adversarial attacks bypassing the watermark through personalized fine-tuning, we introduce the Fidelity Preserving Perturbation Modulation. This approach employs an alternating optimization strategy, iteratively refining both the model parameters and adversarial perturbations, thus attaining an optimal equilibrium between adversarial robustness, image quality, and watermark fidelity.
@@
Our contributions are as follows: (1) We have proposed a novel concept-driven watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. We consider that a refined watermark training and fidelity preserving perturbation are two critical factors in the context of concept watermarking, which offers a fresh perspective on how diffusion model watermarking can be applied in real-world scenarios. (2) We propose a mechanism named Fidelity Preserving Perturbation Modulation for endowing concept watermarks with adversarial properties. Through the alternating optimization of model training and adversarial algorithms, it effectively balances adversarial robustness, image quality, and watermark fidelity. (3) Experiments demonstrate that Conceptwm effectively integrates adversarial perturbations into watermarks and maintains better robustness and accuracy in concept watermark embedding and extraction compared with existing 4 baselines over 6.3% image processing attacks.
2 Related Work
@@
Personalization Techniques for Diffusion Models. Given the substantial computational requirements of training Stable Diffusion from scratch, many methods aim to inject specific concepts into the model through fine-tuning. Currently used fine-tuning techniques include Textual Inversion [10], DreamBooth [27] and LoRA (Low-Rank Adaptation)[14]. Textual Inversion focuses solely on training text embeddings during fine-tuning, allowing concepts to be injected into the text encoder without altering the weights of the U-Net. DreamBooth, on the other hand, fine-tunes the entire U-Net portion of the Stable Diffusion model. Unlike conventional text-to-image fine-tuning, DreamBooth employs prior loss during training to prevent overfitting. LoRA achieves quick and lightweight fine-tuning of Stable Diffusion by training incremental weights in the attention layers of the U-Net.
@@
Model Watermarking for Latent Diffusion Models. WDM [43] trains an autoencoder to stamp a watermark on all training data before re-training the generator from scratch. Aqualora [8] propose a watermark Lora method to watermark all the images. VAE-based methods [36, 17, 9] fine-tune VAE-Decoder to ensure that all generated images contain the watermark. However, the aforementioned watermarking schemes apply watermarks to all generated images, leaving a gap in the protection of specific styles. Meanwhile, A large number of images is often required for the model to learn the watermark embedding pattern, but concept-oriented images for finetuning are often limited.
@@
Protective Perturbation against Stable Diffusion. To safeguard personal images, such as faces and artwork, from potential infringement during the fine-tuning of Stable Diffusion, recent studies focus on disrupting the fine-tuning process by adding imperceptible protective noise to these images. AdvDM [19] introduces a direct adversarial attack on Stable Diffusion by maximizing the Mean Squared Error loss during optimization. This method employs adversarial noise to safeguard personal images. Anti-DreamBooth [31] integrates the DreamBooth fine-tuning process of Stable Diffusion into its framework, designing a bi-level min-max optimization approach to generate protective perturbations. Additionally, other studies [35, 38, 44, 46] have investigated the use of similar adversarial perturbation techniques to generate protective noise for images.
3 Preliminaries
Prompt-based Latent Diffusion Model. Conditional denoising model, denoted as $\epsilon_{\theta}(x,t,c)$, can generate images based on a given text condition $c$. $x_{t}$ represents the latent representation at a particular timestep t within the set $\{1,2...,T\}$. During training, a loss function $\mathcal{L}_{cond}$ is used to guide the diffusion model in denoising the latent representations, $x_{t}=\sqrt{\bar{\alpha}_{t+1}}x_{0}+\sqrt{1-\bar{\alpha}_{t+1}}\epsilon_{\theta}(x_{t})$. Given the input image $x_{0}$, each variable $x_{t}$ is constructed via injecting noise at corresponding timestep $t$. The objective can be expressed as follows:
$\mathcal{L}_{cond}(\theta,x_{0})=\mathbb{E}_{x_{0},t,c,\epsilon\in\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}.$
DreamBooth. DreamBooth employs the base loss of diffusion models in Eq.1. It also incorporates a prior preservation loss to avoid overfitting and text-shifting issues, particularly when only a limited set of instance examples is available. More precisely, it uses a generic prior prompt such as “a photo of [class]” and learn the concept with the prompt such as “a photo of <concept >[class]”. The objective can be expressed as follows:
$\mathcal{L}_{ft}(\theta,x_{0})=\mathbb{E}_{x_{0},t}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}+\lambda\|\epsilon^{\prime}-\epsilon_{\theta}(x_{t+1}^{\prime},t,c_{cls})\|_{2}^{2},$
where $\epsilon$ and $\epsilon^{\prime}$ are both sampled from the standard normal distribution $\mathcal{N}(0,\mathbf{I})$, the noisy variable $x_{t+1}^{\prime}$ is generated from the original stable diffusion model $\theta_{\text{ori}}$ using the prior prompt $c_{\text{cls}}$. The term $\lambda$ emphasizes the importance of the prior term in the generation process.
Adversarial attacks. The goal of adversarial attacks is to find imperceptible perturbations to the input image that mislead the behavior of a given model. Typical research usually focuses on classification tasks, where, for a model $f$, an adversarial example $x^{\prime}$ of an input image $x$ is generated such that it is visually indistinguishable, while causing a misclassification $y=f(x^{\prime})$ The perturbation is typically constrained to lie within an $\eta$-ball to ensure minimal visual difference, specifically satisfying $\|x^{\prime}-x\|_{p}<\eta$. Projected Gradient Descent (PGD) [23] is a widely used attack method that relies on an iterative optimization process. The procedure to update and predict $\tilde{x}$ for an untargeted attack is as follows:
$\tilde{x}^{t+1}=\Pi_{(x,\eta)}\left(\tilde{x}^{t}+\alpha\cdot sign(\nabla_{x}\mathcal{L}(f(x),y))\right),$ (3)
where $\Pi_{x,\eta}(z)$ restricts the pixel values of $z$ within an $\eta$-ball around the original values in $x$.
4 Methodology
Figure 2: Framework of the proposed Conceptwm watermarking scheme for Latent Diffusion Model.
4.1 Motivation
Challenges of existing methods: (1) How to achieve concept-oriented watermark fine-tuning of U-net with limited data and training budget? We will solve this via the framework design in Sec. 4.4. (2) Existing watermarking methods lack protective mechanisms to prevent watermark ”jailbreaking” through personalized fine-tuning. Integrating the existing protective perturbation method into the watermarking mechanism and embedding it into the model while ensuring good generation quality is challenging. This will be solved in Sec. 4.5.
4.2 Overview
The main objective of our method is to protect images of the specific concept generated by the diffusion model by embedding adversarial noise and the watermark pattern into the U-net component of the diffusion model. Figure 2 illustrates an overview of our approach. The Watermark Components Pretraining module implements watermarking in the latent space and enhances the robustness of the watermark through a distortion layer. Fidelity Preserving Perturbation Modulation implements adversarial perturbations to the image while balancing watermarking, image quality, and adversarial properties. Concept-oriented Watermark Fine-tuning enables learning for specific concepts and integrates adversarial watermarks through fine-tuning.
4.3 Watermark Components Pretraining
Existing image watermarking methods often face significant differences between the watermark embedding pattern and the latent space, which causes the watermark information to be disrupted or even lost when transformed into the latent space by the VAE encoder or diffusion regeneration. This issue has been discussed in related watermarking attack methods [1, 3, 5]. Our objective is to embed a robust watermark within the image that is both resilient and conducive to efficient learning by the U-Net architecture.
The U-Net learns the distribution and noise of the image in the latent space, and embedding the watermark into the latent space enables the model to learn the consistency between the watermark and the generated images. In this step, the watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image. The message loss is the Binary Cross Entropy (BCE) between the message $m$ and the sigmoid $\sigma(m^{\prime})$.
$\mathcal{L}_{m}=-\sum_{k=0}^{n-1}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$ (4)
$I_{0}$ is the original image. $I_{r}$ is the reconstructed image. $m$ is secret message (i.e.,watermark). To ensure visual consistency, we use the LPIPS loss function [40] to evaluate the global loss between the watermarked image and the reconstructed image. To avoid excessive modification of local image regions, we select the most prominent parts of the image and calculate channel loss.
$R$ $es(\mathbf{I}_{0},\mathbf{I}_{r})=\frac{1}{3}\sum_{c=1}^{3}|{\mathbf{I}_{c}}^{0}(i,j)-{\mathbf{I}_{c}}^{r}(i,j)|$ (5)
$\mathcal{L}_{channel}=\sum_{k=1}^{K}\left[\operatorname{Top-k}\left(\operatorname{Res}(\mathbf{I}_{0},\mathbf{I}_{r})\right)\right]$
Our training objective can be summarized as follows, where $\lambda$ and $\mu$ represent the coefficients:
$\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{channel}$ (6)
The trained message-encoder and message-decoder will collaborate with the Fidelity Preserving Perturbation Modulation module to assist the U-Net in learning the adversarial watermark.
4.4 Concept-oriented Watermark Fine-tuning
At this stage, the watermark is embedded into specific concepts and then injected into the U-Net. The main challenges are as follows: The fine-tuned images for specific concepts are often limited, making it difficult for the model to learn the specific watermark patterns; Due to the introduction of adversarial perturbations, directly using the DreamBooth training paradigm will be affected by protective perturbations, preventing the model from learning the concept.
The limited number of specific concept samples and adversarial perturbation often result in failures when directly adding an adversarial watermark into images and deploying DreamBooth. This can prevent the watermark embedding pattern from being learned effectively and lead to quality degradation. Our core idea is to decouple concept learning from adversarial watermark learning. In the DreamBooth stage, we learn the specific concept, and by adding adversarial noise to the same image and also using the image without the adversarial noise, the model learns to differentiate between them. The training objective of this stage is as follows:
$\mathcal{L}_{\mathrm{ft}}(\theta)$ $:=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{ori}(x_{t},t,c\big{)}\|^{2}\Big{]}$ (7)
$\varphi$ represents the adversarial perturbation containing the watermark generated through Fidelity Preserving Perturbation Modulation, $\epsilon_{ori}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
4.5 Fidelity Preserving Perturbation Modulation
The goal of this fine-tuning is to maximize the loss of $L_{cond}$ in Eq. 1 and minimize the loss of $L_{ft}$ 2. The adversarial requirement for watermarking is to disrupt the learning process of DreamBooth. Since the DreamBooth model tends to overfit on adversarial images, this can be exploited to make it perform worse when reconstructing clean images.
$\sigma^{*(i)}$ $=\underset{\delta^{(i)}}{\arg\max}\mathcal{L}_{\text{cond }}(\theta^{*},x^{(i)}),\forall i\in\{1,..,N_{ft}\},$ (8)

### v2

# ConceptWM: Watermarking Diffusion Models for Visual Concept Copyright Protection

@@ Abstract.
The personalization techniques of diffusion models succeed in generating specific concepts but also pose threats to copyright protection and illegal use. Model Watermarking is an effective method to prevent the unauthorized use of subject-driven or style-driven image generation, safeguarding concept copyrights. However, under the goal of concept-oriented protection, current watermarking schemes typically add watermarks to all images rather than applying them in a refined manner targeted at specific concepts. Additionally, the personalization techniques of diffusion models can easily remove watermarks. Existing watermarking methods struggle to achieve fine-grained watermark embedding with a few images of specific concept and prevent removal of watermarks through personalized fine-tuning. Therefore, we introduce a novel concept-oriented watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. We introduce Fidelity-preserving Latent Watermarking (FLW) to generate latent watermarks based on image characteristics and the Adversarial Watermarking Modulation module to prevent ”jailbreaking” via personalized finetuning. To enhance U-Net’s efficiency in learning watermark patterns with limited samples, we propose Efficient Concept Watermark Finetuning, which alternates optimization of model parameters for both watermark embedding and concept learning. We conduct extensive experiments and ablation studies to verify our framework. Our code is available at https://anonymous.4open.science/r/Conceptwm-4EB3/.
Latent Diffusion Model, Image Generation, Model Watermarking
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision
1. Introduction
@@
Recent advancements in diffusion models have significantly upgraded image generation techniques, enabling the creation of realistic images from simple text descriptions and meeting the demand for personalized and customized content. The upgraded diffusion models excel in generating novel concept descriptions, such as unique visual themes, motifs, stylistic preferences, or characters, using only a few images of specific themes or styles (Shi et al., 2024; Gal et al., 2023, 2022; Kumari et al., 2023). Through personalization tools, e.g., DreamBooth (Ruiz et al., 2023) and Lora (Hu et al., 2021), any users with minimal design skills can create images of specific concepts by using only a few personal images and simple text inputs. However, the use of personalization techniques introduces potential risks of concept misuse or leakage. Untrustworthy users may exploit these tools to create unauthorized concepts, leading to violations of original creators’ intellectual property and creativity. For instance, creating concepts related to celebrities or artworks can result in deepfake incidents (Juefei-Xu et al., 2022), the spread of fake news and online spam (Barrett et al., 2023), and serious copyright violations.
@@
Figure 1. (a) Post-processing Watermark. (b) U-net Finetuning Watermark. (c) VAE-Finetuning Watermark. (d) Latent-based Watermark. (e) Ours.
@@
Concept watermarking technique is considered to be an alternative to address ownership and accountability issues of specific concepts in generative model, which aims to embed watermark information into the concept and extract the watermarks from subsequent generated images on the basis of the watermarked concept (Feng et al., 2023). In general, the model owner possesses a pre-trained diffusion model and aims to generate images for a new concept. The concept owner provides images of the specific concept to fine-tune the existing diffusion model. During this fine-tuning process, watermark information related to the concept is embedded, enabling the model to generate images with the new concept. A watermark decoder can extract the watermark information from the generated images.
@@
However, as shown in Fig. 1, existing watermarking methods have not yet to achieve concept watermarking, despite the growing interest in adding watermarks to generated images (Kim et al., 2024; Xiong et al., 2023; Zhang et al., 2024; Yang et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b; Wen et al., 2024; Feng et al., 2024; Cox et al., 2007). On the one hand, prior work has rarely addressed fine-grained watermarking mechanisms for specific concepts. Most existing methods (Wen et al., 2024; Fernandez et al., 2023; Zhao et al., 2023b) watermark the entire images rather than images of specific concepts, making them inapplicable for concept watermarking scenario. A major challenge in concept watermarking is to ensure that the watermark not only exists within a specific image but also remains present across various image variations genraeted by the diffusion model. Compared to traditional image watermarks embedded at the pixel level, concept watermarks appears a higher-level technical barrier, as it associates the watermarks with high-level sematic features related to specific concepts. Since the number of images with a specific concept is limited, embedding watermarks in personalization models is a great challenge, especially when fine-tuning the model by injecting a new concept.
@@
On the other hand, robustness remains a challenge for current watermarking methods in concept protection scenarios, when considering diverse adversaries. Recent studies have demonstrated that existing methods are ineffective against model purification attacks (Zhao et al., 2024; Liu et al., 2024), as adversaries can bypass watermarking schemes by generating watermark-free images through personalized fine-tuning. The technical challenge lies in balancing adversariality, image quality, and watermark fidelity. One potential approach to enhancing watermark adversariality is to inject imperceptible protective adversarial perturbations into images, even though this type of methods frequently compromises both image quality and watermark fidelity (Liang et al., 2023; Shan et al., 2023; Van Le et al., 2023). Therefore, there is an urgent need for watermarking solution that ensures concept tracing while effectively addressing adversarial threats.
@@
Our investigation finds three fundamental properties required by model watermarking, including refined nature, robustness, and adversarial properties. To be specific, refined nature, as a core component of diffusion model, requires the ability to embed a watermark into a specific concept through minimal training data and steps without compromising fidelity, as the knowledge of a specific concept generally is learnt by the denoising network (U-Net). Robustness ensures that the watermark remains stable across various image processing operations and generation parameters. Adversarial properties typically guarantee that the generated concept-specific images cannot be easily “jailbreaken” via personalized fine-tuning.
@@
To address these challenges, we propose a novel concept watermarking framework, named ConceptWM, which focuses on two critical factors, namely refined watermark training and adversarial perturbation modulation. To achieve the concept watermark embedding, we develop a Efficient Concept Watermark Finetuning (ECWF) scheme, which rapidly learns watermark embedding patterns in diffusion models with limited data and training resources. Additionally, we develop a Fidelity-preserving Latent Watermark (FLW) module that incorporates the watermark into the latent space and employs a distortion layer in order to enhance robustness against image processing attacks. Next, we propose an Adversarial Watermark Modulation (AWM) to balance adversarial perturbation, watermark fidelity, and image quality, while mitigating the risk of adversarial attacks bypassing the watermark through personalized fine-tuning. This scheme utilizes an alternating optimization strategy to refine both referenced model parameters and adversarial perturbations, so that an optimal equilibrium among adversarial robustness, image quality, and watermark fidelity can be achieved.
Figure 2. Framework of the proposed ConceptWM watermarking scheme for Latent Diffusion Model. The Efficient Concept Watermark Finetuning module employs a two-step training process to learn specific concepts and watermark embedding patterns. The Adversarial Watermark Modulation module utilizes a surrogate model fully fine-tuned on a referenced clean image set $X_{B}$ to guide PGD optimization, iteratively generating adversarial watermarks through multiple rounds of watermark and noise integration.
@@
Our contributions are as follows: (1) We propose a novel concept-driven watermarking framework that seamlessly embeds imperceptible watermarks into the concept of diffusion models. The proposed framework highlights the importance of refined watermark training and fidelity-preserving perturbations in real-world applications of diffusion model watermarking. (2) We propose a mechanism named Adversarial Watermark Modulation that endows concept watermarks with adversarial properties. Through the alternating optimization of model training and adversarial algorithms, it effectively balances adversarial robustness, image quality, and watermark fidelity. (3) Experiments demonstrate that ConceptWM effectively integrates adversarial perturbations into watermarks and maintains better robustness and accuracy in concept watermark embedding and extraction compared with existing 4 baselines over 6.3% image processing attacks.
2. Related Work
@@
Personalization Techniques for Diffusion Models. Recent explorations mainly focus on injecting specific concepts into the model through fine-tunning tools, e.g., Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023) and LoRA (Low-Rank Adaptation)(Hu et al., 2021), given the substaintial computational requirements of training Stable Diffusion from scratch. Specifically, Textual Inversion focuses on training text embedding so that concepts are allowed to be injected into the text encoder without altering weights of the U-Net. Dissimilar to conventioanl text-to-image fine-tuning, DreamBooth fine-tunes the entire U-Net portion of the Stable Diffusion model and employs prior loss during training to prevent overfitting. LoRA is designed to achieve lightweigth efficient fine-tuning by training incremental weights in the attention layers of the U-Net. Our work focuses on watermarking specific concepts generated through DreamBooth fine-tuning.
@@
Model Watermarking for Latent Diffusion Models. The embedding characteristics of different watermarking methods for latent diffusion models are illustrated in Figure 1. It is mainly divided into Post-processing Watermark, U-Net Finetuning Watermark, VAE-Finetuning Watermark, and Latent-based Watermark. These watermarking methods aim to protect the copyright of images or generative models, so they often embed watermarks into all images. For example, WDM (Zhao et al., 2023b) is designed to train an autoencoder to stamp a watermark on all training data before re-training the generator from scratch . Aqualora (Feng et al., 2024) is a lora-based method that watermarks all images. VAE-based methods (Xiong et al., 2023; Kim et al., 2024; Fernandez et al., 2023) fine-tune VAE-Decoder to ensure all generated images containing the watermark. However, this all-image type of methods encounters challenges in protecting specific styles. A large volume of images are generally required for the model to learn the watermark embedding pattern, but concept-oriented images for finetuning are limited to only a few.
@@
Protective Perturbation against Stable Diffusion. Recent studies mostly have addressed disrupting the fine-tuning process by adding imperceptible protective noise to images, in order to prevent personal images from potential infringement during the fine-tuning of Stable Diffusion. For example, AdvDM employs adversarial noise to safeguard personal images on Stable Diffusion by maximizing the Mean Squared Error loss during optimization (Liang et al., 2023). Anti-DreamBooth (Van Le et al., 2023) offers a bi-level min-max optimization for generating protective perturbations, through integrating the DreamBooth fine-tuning process of Stable Diffusion into its framework. Other studies (Wu et al., 2023; Ye et al., 2023; Zhao et al., 2023a; Zheng et al., 2023) have explored similar adversarial perturbation techniques to generate protective noise for images.
3. Methodology
3.1. Overview
The main objective of our scheme is to protect images of a specific concept that are generated by the diffusion model, by embedding an adversarial watermark pattern into the U-Net component of the diffusion model. Fig. 2 illustrates an overview of our approach. To embed a watermark that is both fidelity-preserving and conducive to learning by U-Net, we develop a Fidelity-preserving Latent Watermarking (FLW) module, which adds a combination layer for the message encoder and latent concatenation, so that watermarks can be generated based on image characteristics rather than a fixed watermark signal. To protect the generated concept-specific images from being “jailbroken” through personalized finetuning, we develop an Adversarial Watermarking Modulation module to integrate adversarial propeties into the watermark. To enable U-Net to efficiently learn watermark patterns during concept learning with limited samples, we present an Efficient Concept Watermark Finetuning module that alternates optimization of model parameters to simultaneously achieve adversarial watermark embedding and concept learning.
3.2. Fidelity-preserving Latent Watermark
Figure 3. The comparison of two variants of Latent watermark namely Overlay-based Latent Watermarking and Fusion-based Latent Watermarking.
Existing image watermarking techniques frequently arise substantial discrepancies between the watermark embedding pattern and the latent space. This mismatch issue may cause the disruption or even complete loss of watermark information when the image is encoded into the latent space by a VAE encoder or regenerated through diffusion, as has been discussed by recent watermarking attacks (An et al., 2024; Ballé et al., 2018; Cheng et al., 2020). Thus, we aim to embed a robust watermark within the image that is both resilient and conducive to efficient learning for U-Net. To be specific, we have developed two variants of latent watermarking, which are Overlay-based Latent Watermarking (OLW) and Fidelity-preserving Latent Watermarking (FLW). As shown in Fig. 3, both variants involve the latent domain perturbations. The watermarked images are processed through the message decoder to evaluate the discrepancy between the decoded watermark information and the original content of the watermarked image. The message loss is the Binary Cross Entropy (BCE) between the message $m$ and $\sigma(m^{\prime})$. Given the message length $N$, the message loss $\mathcal{L}_{m}$ is defined in Eq. (1).
(1) $\mathcal{L}_{m}=-\sum_{k=1}^{N}m_{k}\log\sigma(m_{k}^{\prime})+(1-m_{k})\log(1-\sigma(m_{k}^{\prime}))$
Overlay-based Latent Watermarking. OLW essentially overlays a fixed watermark signal that can be removed through subtraction. We observe that OLW-watermarked images may face the issue of the watermark integrity, since noticeable artifacts appear in local regions of the images, making the images distinguishable to the huaman eye.
Fidelity-preserving Latent Watermarking. To address issues of OLW, we have developed Fidelity-preserving Latent Watermarking (FLW) that adds a combination layer after the message encoder and latent concatenation. To ensure visual consistency, we calculate LPIPS loss $\mathcal{L}_{pips}$ (Zhang et al., 2018) and utilize a Peak Pixel Difference (PPD) Loss $\mathcal{L}_{ppd}$ which selects the most prominent parts of the image and calculates channel loss to reduce excessive modification of the images. Our training objective ($\mathcal{L}_{total}$) can be expressed in Eq. (2), where $\lambda$ and $\mu$ represent the coefficients.
(2) $\mathcal{L}_{total}=\mathcal{L}_{m}+\lambda\mathcal{L}_{pips}+\mu\mathcal{L}_{ppd}$
The jointly optimized message encoder and decoder, working in concert with the Adversarial Watermark module, will empower the U-Net to learn robust watermarking patterns.
3.3. Adversarial Watermark Modulation
To protect the generated concept-specific images from “jailbreaking” through personalized fine-tuning, we integrate adversarial properties into the watermark. By maximizing the image generation loss with adversarial noise added to the watermark and minimizing the concept learning loss through Dreambooth fine-tuning with reference images, we ensure that the adversarial watermark does not interfere with concept learning while preventing “jailbreaking” through personalized fine-tuning.
Image Generation Loss. To enable the conditional denoising model $\epsilon_{\theta}(x,t,c)$ to match the data distribution of the target images, $\mathcal{L}_{cond}$ is a loss function that ensures the conditional denoising model generates images conditioned on a given text condition ($c$). Given the input image $x_{0}$, the $\mathcal{L}_{cond}$ is as follows in Eq. (3):
$\mathcal{L}_{cond}(\theta,x_{0})=\mathbb{E}_{x_{0},t,c,\epsilon\in\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}$
where each variable $x_{t}$ is constructed through noise injection at a corresponding timestep $t$ within the set $\{1,2,\cdots,T\}$.